# Layer Swapping for Zero-Shot Cross-Lingual Transfer in Large Language Models

### v1

@@ 5.1 Experiments on Swahili, Telugu, Japanese, and Bengali
As displayed in Table 1, we find that layer swapping consistently outperforms these baselines on MGSM. For Swahili and Telugu, the default layer swapping configuration scores higher than both the individual experts across all 9 pairs. This provides tremendous evidence that this merging method prevents negative interference. For Bengali, where the math expert already performs significantly better than for Swahili and Telugu, numerous layer-swapping configurations consistently surpass the math expert’s average score. In addition, the best layer swapped model performs higher than the best math expert on its own. For Japanese, we find here lower average performance compared to the individual math experts. However, we note that our Japanese experts were perhaps the weakest; the increase in performance across Belebele, FLoRes, MBPP, and MMLU after SFT were minimal and with the data and hyperparameters tested, we were unable to do better. Furthermore, MGSM performance of both the base Llama 3.1 and the math experts were already decent in Japanese, prior to merging. In comparison to model souping, layer swapping consistently outperforms it in terms of both maximum and average performance in all languages. Model soup results typically lie in between the individual results of the experts.
We further evaluated these layer-swapped models on Belebele, FLoRes, MBPP, and MMLU to ensure no over-optimization to this benchmark and find that the results are on par, usually even higher, than the base model and the individual experts. In addition, we manually checked a small number of text generations to further ensure the quality of the resulting model. In addition, we also evaluate combining layer swapping with model souping given that there are several language and math experts. We experiment souping with souping the three experts for each of the languages before layer swapping with a soup of the three math experts and present results in Table 5 in the Appendix. We find that these language and math soups perform on par with the individual pairs when layer swapped together. This feasibility in this settings shows that layer swapping can further extend the search space of fine-tuned models.
Given these results, we conjecture that perhaps layer swapping provides the biggest benefit for lower-resource languages (e.g. Swahili, Telugu), although such a conclusion would require evaluating more than four languages. This conjecture would be explained by the difficulty to improve language capabilities with low-data SFT when Llama 3.1 was already pretrained on more tokens in that language. Additionally, for higher-resource languages, cross-lingual transfer occurs more naturally when fine-tuning on English math data.
@@ 5.2 Further Evaluations on Swahili
Given that here,model souping appears to be susceptible to negative interference when merging, we additionally evaluate TIES-merging (Yadav et al., 2023) for Swahili. This merging method resolves interference between experts by sequentially trimming conflicting values, determining the sign of each weight, and then merging the resulting weights. In addition, we present an alternative for Swahili where instead of post hoc model merging, we mix the language and math datasets and do a single fine-tuning run with samples from both. Instead of training two experts, each on 30-40k samples, the joint SFT is over the union (80k samples). Identical to the training of experts, we launch many SFT runs with different hyperparameters and select the best three checkpoints.
@@
Table 2 displays these more detailed results for Swahili. We find that TIES-merging was similarly performing as model souping in our setup, which means consistently worse than layer swapping. For the Swahili and math data mixing prior to fine-tuning, we are able to achieve comparable results to layer swapping on average, with the average MGSM score for this joint SFT being less than one point lower. However, the maximum performance of these checkpoints lagged the best layer-swapped pairs by 4.4 percentage points. This means that the ceiling for cross-lingual transfer is significantly higher with this methodology than simply mixing datasets together. This is significant because our method of merging two variants fine-tuned on separate datasets proves more effective than an extended fine-tuning on the combined datasets.
We note that by swapping in layers from the Swahili expert, MGSM performance in English and other non-Swahili languages decreases from the math expert. This is expected given we optimize for a different language, yet the decrease is relatively small nonetheless. For comparison, we see that performance drops much more for the other merging methods on these non-target languages.
@@ 6 Discussion
Layer swapping is a simple, yet effective method for merging together models for cross-lingual transfer. The success of this post hoc method prompts numerous key insights.
@@
In settings where in-language task data is unavailable or rare, such as labeled math samples in Telugu, layer swapping provides a very effective way to create a capable model with simply English task data and general target language data. For lower-resource languages, such a constrained scenario is extremely common in practice—especially when it comes to instruction fine-tuning data. All other baselines we evaluated in such a scenario were not able to combine language and math capabilities as consistently and as effectively, as discussed in Section 5.
Model souping is popular because of its flexibility, convenience, and ability to expand model search space. In this regard, layer swapping provides a more effective alternative for the multilingual setting with the same advantages. This method can be implemented fully post hoc between any number of checkpoints using simple parameter arithmetic, with no further fine-tuning. Because of how inexpensive it is to implement, it enables the ability to quickly iterate through and test many configurations (e.g. the number of swapped transformer layers). Similar to souping, the simplicity of this method means that it has the potential to work at any stage of training (pretraining, fine-tuning, preference tuning, etc.). In multi-task fine-tuning scenarios, this could allow for non-English capabilities being boosted on the side, separately from other reasoning capabilities—such as math, multi-hop reasoning, coding, safety, etc.—and then models are merged back together post hoc to combine the learned skills.
@@
The effectiveness of this method indicates that there are meaningful multilingual patterns occurring in these LLMs that researchers can uncover and exploit. Potentially, we can reinterpret the functionality of the top and bottom transformer layers as, to a certain degree, natural language interfaces to broader model intelligence. Accordingly, this can potentially give way to a newer generation of modular solutions for multilingualism in decoder-only LLMs, similar to the MoE or adapter solutions that were designed for encoder models (Pfeiffer et al., 2020; 2022; Liu et al., 2023). Regardless, this work suggests that further interpretability research into LLM multilinguality could lead to more cost-effective solutions for boosting non-English capabilities.
@@
In terms of model merging, our analysis indicates that weight-level techniques are perhaps not as effective for reducing interference when merging parameters together, as compared to parameter-level or even layer-level merges. A possible explanation is that modifying individual weight values may disrupt the linear dependence within the transformations defined by the weight tensors. Conversely, the recent success of self-speculative decoding in LLMs (Zhang et al., 2024a; Elhoushi et al., 2024) suggests that the output embedding space of most layers across the model are very similar. If so, keeping entire layers in tact may alleviate the risk of undermining newfound learnings. Another potential reason is that using very granular magnitude measures may be too noisy to properly indicate which weight updates are important and which are not. By the nature of backpropagation, weight magnitude changes may be more indicative of the magnitude of the original value (larger weights will have larger gradients) or simply the weight’s proximity to the output layer. This is perhaps especially the case for small-scale training runs, where there is more noise.
@@ 7 Limitations and Future Work
Freezing parameters before training experts: Instead of re-composing a model with layers from separate models, a logical next evaluation would be to freeze model layers from the beginning. Intuitively, this would only be more effective from a performance standpoint as there is no ad hoc merging of experts from disparate training jobs. That being said, part of the benefit of this solution is that it is convenient and flexible because it’s all post hoc and required no prior planning. Parameter freezing would require knowing which configuration would work prior to training.
LLMs with different pretraining: Any transformer model should work similarly, but it is unclear if the English-centric nature of Llama 3.1 enabled the success of layer swapping (or conversely inhibited it). It is also possible that in weaker models (e.g. Llama 1), the non-English representations are not developed enough to enable this cheap transfer. Similarly, our analysis may not extend to LLMs with more balanced multilingual pretraining, who may have very different representation spaces.
Different model sizes: LLMs with more parameters, from either more transformer layers or larger hidden dimensions, would clearly require new merging configurations. In addition, it remains to be seen how related attributes such as sparsity, redundancy, and the quantity of cross-lingual shared parameters, would influence the performance of layer swapping.
@@
Other reasoning tasks: This work only tackles mathematical reasoning, but it’s conceivable that other complex reasoning capabilities are concentrated in parameters separate from language capabilities in a similar manner. That being said, the configuration required for layer swapping may be completely different and may require a preliminary analysis as in Section 3.
Parameter-efficient fine-tuning: Since layer swapping treats transformer layers as unified blocks, it would be equivalent if the model were fine-tuned using parameter-efficient SFT methods such as LoRA, which consists of inserting adapter weights into the transformer layer (Hu et al., 2022). The modification of other model merging methods for such adapters is simple (Zhang et al., 2023a), and therefore layer swapping has the potential to be effective in parameter-efficient fine-tuning settings.
Different training stage: We limit the focus of this work on low-data SFT. Model souping, itself, is implemented in practice at all stages (e.g. pretraining, CPT). This raises the question of in which settings layer swapping would be effective for model merging. Similarly, model souping is effective even when the checkpoints have been fine-tuned significantly. It is unclear what would be the point for our methodology where the experts have diverged from each other too significantly that they would no longer recombine well. If it is only effective in small training runs, it could still enable a multi-step training where experts are iteratively fine-tuned and merged successively, analogous to gradient aggregation in data-distributed training.
@@ 8 Conclusion
In this paper, we address the difficulty of training LLMs to perform tasks in languages where labeled task-specific data does not exist. We first trained separate experts on English math data and generic instruction data in numerous target languages. An analysis of the importance of different parameter updates during fine-tuning led to the development of the layer swapping method which swaps in the top and bottom layers from language experts into the math expert. Surprisingly, this simple model merging provides one of the highest performing methods to fine-tune an LLM for math in a target language without the presence of target language data. The strong intuition behind this solution and its flexibility, simplicity, and effectiveness provides vast potential for being practical in many other settings. This method can be potentially adapted for new pretrained models, target tasks, target languages, training stages, training setups, and more. In addition, this method indicates better model interpretability of multilinguality can lead to more efficient methods for transferring English capabilities to lower resource languages.
@@
For reproducing the fine-tuning of our “expert” models, we describe the process in Section 3.1 and provide further details on the data and fine-tuning hyperparameters in Appendix A.1 and A.2. Further details for our analysis described in Section 3.2 can be found in Appendix A.4. We provide pseudocode of the layer swapping algorithm defined in Section 4 in Algorithm 1. Our experiments for which results are presented are all explained thoroughly in Section 5.
@@ A.2 Supervised Fine-tuning Details
Expert Learn Rate Batch Size Seq. Length weight decay clip, max norm sched. warmup $\beta 2$
@@
A.3 Results from combining with souping
@@
A.4 Parameter-Level Visualizations of the Expert Vectors
Larger parameter-level visualizations for a number of the experts, each visualization is the same as described in Fig. 2

### v2

@@ 5.1 Experiments on Swahili, Telugu, Japanese, and Bengali
As displayed in Table 1, we find that layer swapping consistently outperforms these baselines on MGSM. For Swahili and Telugu, the default layer swapping configuration scores higher than both the individual experts across all 9 pairs. This provides tremendous evidence that this merging method prevents negative interference. For Bengali, where the math expert already performs significantly better than for Swahili and Telugu, numerous layer-swapping configurations consistently surpass the math expert’s average score. In addition, the best layer swapped model performs higher than the best math expert on its own. For Japanese, here we find a lower average performance compared to the individual math experts. However, we note that our Japanese experts were perhaps the weakest; the increases in performance across Belebele, FLoRes, MBPP, and MMLU after SFT were minimal and with the data and hyperparameters tested, we were unable to do better. Furthermore, MGSM performance of both the base Llama 3.1 and the math experts was already decent in Japanese, prior to merging. Compared to model souping, layer swapping consistently outperforms it in terms of both maximum and average performance in all languages. The model soup results typically lie between the individual results of the experts.
We further evaluated these layer-swapped models on Belebele, FLoRes, MBPP, and MMLU to ensure no over-optimization to the MGSM benchmark and find that the results are on par, usually even higher, than the base model and the individual experts. In addition, we manually checked a sample of text generations to further ensure the quality of the resulting model. Given that there are several language and math experts, we evaluate combining layer swapping with model souping. We do so by souping the three experts for each of the languages before layer swapping with a soup of the three math experts and present results in Table 6 in the Appendix. We find that these language and math soups perform on par with the individual pairs when layer swapped together. The feasibility in these settings shows that layer swapping can further extend the search space for fine-tuned models.
Given these results, we conjecture that perhaps layer swapping provides the biggest benefit for lower-resource languages (e.g. Swahili, Telugu), although such a conclusion would require evaluating more than four languages. This conjecture would be explained by the difficulty of improving language capabilities with low-data SFT when Llama 3.1 was already pretrained on more tokens in that language. Additionally, for higher-resource languages, cross-lingual transfer occurs more naturally when fine-tuning on English math data.
@@ 5.2 Further Evaluations on Swahili
Given that here, model souping appears to be susceptible to negative interference when merging, we additionally evaluate TIES-merging (Yadav et al., 2023) for Swahili. This merging method resolves interference between experts by sequentially trimming conflicting values, determining the sign of each weight, and then merging the resulting weights. In addition, we present an alternative for Swahili where, instead of post hoc model merging, we mix the language and math datasets and do a single fine-tuning run with samples from both. Instead of training two experts, each on 30-40k samples, the joint SFT is over the union (80k samples). Identical to the training of experts, we launch many SFT runs with different hyperparameters and select the best three checkpoints.
@@
Table 2 displays these more detailed results for Swahili. We find that TIES-merging, similar to model souping, consistently underperforms compared to layer swapping. For mixing the Swahili and math data prior to fine-tuning, we are able to achieve results comparable to layer swapping on average, with the average MGSM score for this joint SFT being less than one point lower. However, the maximum performance of these checkpoints lagged the best layer-swapped pairs by 4.4 percentage points. This means that the ceiling for cross-lingual transfer is significantly higher with this methodology than simply mixing datasets together. This is significant because our method of merging two variants fine-tuned on separate datasets proves to be more effective than an extended fine-tuning on the combined datasets, consistent with Aakanksha et al. (2024).
We note that by swapping in layers from the Swahili expert, MGSM performance in English and other non-Swahili languages decreases from the math expert. This is expected given that we optimize for a different language, yet the decrease is relatively small nevertheless. For comparison, we see that performance drops much more for the other merging methods on these non-target languages.
@@ 6 Discussion
Layer swapping is a simple, yet effective method for merging models for cross-lingual transfer. The success of this post hoc method prompts numerous key insights.
@@
In settings where in-language task data is unavailable or rare, such as labeled math samples in Telugu, layer swapping provides a very effective way to create a capable model with simply English task data and general target language data. For lower-resource languages, such a constrained scenario is extremely common in practice, especially when it comes to instruction fine-tuning data. All other baselines that we evaluate in such a scenario are not able to combine language and math capabilities as consistently and effectively, as discussed in Section 5.
@@
The surprising effectiveness of this method can be explained by the observation that English-centric LLMs do most of their internal thinking in English, which aligns with recent literature (Wendler et al., 2024; Zhang et al., 2024b; Tang et al., 2024; Kojima et al., 2024). Therefore, our method exploits the finding that the most important language-specific parameters are those in the top and bottom transformer layers that map multilingual input to and from English representations. Potentially, we can reinterpret the functionality of these transformer layers as, to a certain degree, natural language interfaces to broader model intelligence. Our method preserves the updates to the middle layers, as we hypothesize that this is where most of the improvements to the LLM’s mathematical reasoning are concentrated. Consequently, this separability of multilingual parameters from other model capabilities could motivate a new generation of modular solutions for multilingualism in decoder-only LLMs—akin to MoE or adapter-based approaches designed for encoder models (Pfeiffer et al., 2020; 2022; Liu et al., 2023). Regardless, this work suggests that further interpretability research into LLM multilinguality could lead to more cost-effective solutions to boost non-English capabilities.
Model souping is popular because of its flexibility, convenience, and ability to expand the model search space. In this regard, layer swapping provides a more effective alternative for the multilingual setting with the same advantages. This method can be implemented fully post hoc between any number of checkpoints using simple parameter arithmetic, with no further fine-tuning. Because of how inexpensive it is to implement, it enables the ability to quickly iterate through and test many configurations (e.g. the number of swapped transformer layers). Similarly to souping, the simplicity of this method means that it has the potential to work at any stage of training (pretraining, fine-tuning, preference tuning, etc.). In multi-task fine-tuning scenarios, this could allow for non-English capabilities being boosted on the side, separately from other reasoning capabilities—such as math, multi-hop reasoning, coding, safety, etc.—and then models are merged back together post hoc to combine the learned skills.
@@
In terms of model merging, our analysis indicates that weight-level techniques are perhaps not as effective in reducing interference when merging parameters together, as compared to parameter-level or even layer-level merges. A possible explanation is that modifying individual weight values may disrupt the linear dependence within the transformations defined by the weight tensors. Conversely, the recent success of self-speculative decoding in LLMs (Zhang et al., 2024a; Elhoushi et al., 2024) suggests that the token representation spaces of most layers across the model are very similar. If so, keeping entire layers in tact may alleviate the risk of undermining newfound learning. Another reason could be that using very granular magnitude measures may be too noisy to properly indicate which weight updates are important and which are not.
@@ 7 Limitations and Future Work
Freezing parameters before training experts: Instead of re-composing a model with layers from separate models, a logical next evaluation would be to freeze model layers from the beginning. Intuitively, this would only be more effective from a performance standpoint, as there is no ad hoc merging of experts from disparate training jobs. However, a significant benefit of our solution is that it is all post hoc and flexible. Parameter freezing would require knowing which configuration would work prior to training or require numerous training runs to find the best configuration.
LLMs with different pretraining: An explanation for why this method works so well is that English-centric LLMs “think in English”, which is shown in more than Llama 3.1. It must be studied whether LLMs with balanced multilingual pretraining have similarly consolidated representations, and if not, whether layer swapping breaks.
Different model sizes: LLMs with more parameters, either from more transformer layers or from larger hidden dimensions, would clearly require new merging configurations. In addition, it remains to be seen how related attributes such as sparsity, redundancy, and the quantity of cross-lingual shared parameters, would influence the performance of layer swapping.
@@
Other reasoning tasks: This work only tackles mathematical reasoning, but it is conceivable that other complex reasoning capabilities are concentrated in parameters separate from language capabilities in a similar manner. That being said, the configuration required for layer swapping may be completely different and may require a preliminary analysis as in Section 3.
Parameter-efficient fine-tuning: Since layer swapping treats transformer layers as unified blocks, it would be equivalent if the model were fine-tuned using parameter-efficient SFT methods such as LoRA, which consists of inserting adapter weights into the transformer layer (Hu et al., 2022). Modifying other model merging methods for such adapters is simple (Zhang et al., 2023), and therefore layer swapping has the potential to be effective in parameter-efficient fine-tuning settings.
Different training stage: We limit the focus of this work on low-data SFT. Model souping, itself, is implemented in practice at all stages (e.g. pretraining, CPT), and layer swapping has the potential to be effective there as well. Similarly, model souping is effective even when the checkpoints have been fine-tuned significantly. It is unclear what would be the point for our methodology where the experts have diverged too significantly that they would no longer recombine well. Either way, it could still enable a multi-step training where experts are iteratively fine-tuned and merged successively, analogous to gradient aggregation in data-distributed training.
@@ 8 Conclusion
In this paper, we address the difficulty of training LLMs to perform tasks in languages where labeled task-specific data does not exist. We first trained separate experts on English math data and generic instruction data in numerous target languages. An analysis of the importance of different parameter updates during fine-tuning led to the development of the layer swapping method which swaps in the top and bottom layers from language experts into the math expert. Surprisingly, this simple model merging provides one of the highest performing methods to fine-tune an LLM for math in a target language without the presence of target language data. The strong intuition behind this solution and its flexibility, simplicity, and effectiveness provide vast potential to be practical in many other settings. This method can be potentially adapted for new pretrained models, target tasks, target languages, training stages, training setups, and more. In addition, this method indicates better model interpretability of multilinguality can lead to more efficient methods for transferring English capabilities to lower resource languages.
@@
To reproduce the fine-tuning of our “expert” models, we describe the process in Section 3.1 and provide further details on the data and fine-tuning hyperparameters in Appendix A.1 and A.2. Further details for our analysis described in Section 3.2 can be found in Appendix A.5. We provide pseudocode of the layer swapping algorithm defined in Section 4 in Algorithm 1. All of our experiments for which results are presented are thoroughly explained in Section 5.
@@ A.2 Supervised Fine-tuning Details
Expert Learn Rate Batch Size Seq. Length weight decay clip, max norm sched. warmup steps $\beta 2$
@@
A.3 Performance of individual experts
Table 5: Comprehensive results for each of the set of experts on tasks in the target language. Each number represents an average over the 3 selected experts. The “10-lang avg” column presents the results across 10 languages (English, German, French, simplified Chinese, Russian, Spanish, Japanese, Bengali, Swahili, Telugu).
Task Language Llama 3.1 8b Average across 3-experts
Math Japanese Bengali Swahili Telugu
MGSM English 0.560 0.663 0.561 0.563 0.557 0.561
em@maj1, 8-shot Japanese 0.336 0.427 0.359
Bengali 0.292 0.383 0.335
Swahili 0.248 0.295 0.247
Telugu 0.120 0.201 0.200
10-lang avg 0.311 0.454
Belebele results English 0.871 0.877
accuracy, 5-shot Japanese 0.760 0.754 0.766
Bengali 0.649 0.654 0.683
Swahili 0.553 0.557 0.584
Telugu 0.546 0.557 0.602
FLoRes results Eng-Jpn 20.6 20.5 20.5
BLEU, 0-shot Jpn-Eng 28.5 28.5 28.9
Eng-Ben 18.8 18.8 17.9
Ben-Eng 30.5 29.7 29.0
Eng-Swh 12.5 12.9 13.0
Swh-Eng 31.5 31.5 30.2
Eng-Tel 15.7 15.2 16.5
Tel-Eng 29.2 29.0 27.6
MBPP result Japanese 0.464 0.470 0.462
pass @ 1, 0-shot Bengali 0.438 0.436 0.462
Swahili 0.402 0.407 0.407
Telugu 0.410 0.412 0.437
MMLU result Japanese 0.520 0.518 0.518
accuracy, 5-shot Bengali 0.422 0.420 0.436
Swahili 0.419 0.420 0.425
Telugu 0.415 0.412 0.427
A.4 Results from combining with souping
@@
A.5 Parameter-Level Visualizations of the Expert Vectors
Larger parameter-level visualizations for a number of the experts, each visualization is the same as described in Figure 2